### v2

# GitBugs: Bug Reports for Duplicate Detection, Retrieval Augmented Generation, and More

Abstract
Bug reports provide critical insights into software quality, yet existing datasets often suffer from limited scope, outdated content, or insufficient metadata for machine learning. To address these limitations, we present GitBugs—a comprehensive and up-to-date dataset comprising over 150,000 bug reports from nine actively maintained open-source projects, including Firefox, Cassandra, and VS Code. GitBugs aggregates data from Github, Bugzilla and Jira issue trackers, offering standardized categorical fields for classification tasks and predefined train/test splits for duplicate bug detection. In addition, it includes exploratory analysis notebooks and detailed project-level statistics, such as duplicate rates and resolution times. GitBugs supports various software engineering research tasks, including duplicate detection, retrieval augmented generation, resolution prediction, automated triaging, and temporal analysis. The openly licensed dataset provides a valuable cross-project resource for benchmarking and advancing automated bug report analysis. Access the data and code at https://github.com/av9ash/gitbugs/.
I Introduction
Bug reports are foundational to software maintenance, reliability, and quality assurance. They capture critical user feedback and developer annotations that help identify defects, propose fixes, and prioritize engineering resources. Despite their importance, publicly available datasets of bug reports often fall short in scale, consistency, and breadth. Many existing resources focus on single projects, lack standardized metadata, or are outdated.
To address these challenges, we introduce GitBugs—a comprehensive, curated dataset containing over 150,000 bug reports spanning nine major open-source projects. We designed this dataset to support the software engineering research community by providing clean, labeled, and diverse bug tracking data. GitBugs enables empirical investigations and development of automated techniques across domains such as duplicate detection, bug triaging, resolution classification, and retrieval augmented generation.
Our key contributions are as follows:
  • A large-scale dataset of over 150,000 bug reports collected from 9 open-source software projects, covering Github, Jira and Bugzilla ecosystems.
  • Comprehensive metadata per report, including fields such as Summary, Description, Status, Priority, Resolution, and timestamp information.
  • Project-wise analytics including duplicate rates, time-to-resolution distributions, and metadata completeness.
  • Reproducible artifacts: Duplicate Bug reports mappings, exploratory data analysis (EDA) notebooks, model training and, validation scripts.
We intend this dataset to serve as a benchmark for research and development in bug report understanding and automation, and a foundation for training large-scale models on realistic, task-specific software engineering data.
II Related Work
The software engineering community has long recognized the value of curated bug report datasets in advancing research areas such as defect prediction, localization, duplicate detection, and reproducibility. Over the years, several influential datasets have been released, each tailored to specific research objectives.
II-A Existing Datasets
One of the earliest large-scale efforts in bug report analysis was by Lamkanfi et al. [1], who published a dataset comprising over 200,000 bug reports from Eclipse and Mozilla. Their work highlighted the evolution of bug discussions, laying a foundation for bug triaging and prioritization research.
To support duplicate bug report detection, the LogPAI team introduced BugRepo [2], a collection of free-text bug reports from projects such as Firefox and Eclipse JDT. BugRepo has become a benchmark for applying NLP techniques to software engineering tasks by focusing on textual similarity and clustering of duplicates.
Reproducibility has emerged as another core challenge in bug tracking. In response, Wendland et al. [3] released AndroR2, a dataset of 90 Android bugs accompanied by reproduction steps, scripts, and APKs. This resource has facilitated research on automated bug reporting and empirical reproducibility studies.
Other datasets emphasize code-level analysis. Ferenc et al. [4] introduced BugHunter, which links buggy and fixed code segments with a comprehensive set of code metrics, enabling deeper investigation into defect prediction.
Cross-language bug localization has also gained traction. Muvva et al. [5] developed BuGL, a multilingual dataset containing over 10,000 bug reports from C, C++, Java, and Python projects, aimed at evaluating language-agnostic localization models.
Recent large-scale mining efforts have further enriched the landscape. Oliva et al. [6] curated a longitudinal dataset mapping bug reports to fix commits across 55 Apache projects, providing insights into long-term bug lifecycles. Complementing this, RegMiner [7] automatically collects reproducible regression bugs from 66 projects, supporting research in regression detection and automated repair.
Several other widely-used datasets also exist. The PROMISE repository [8] provides datasets for defect prediction, while Defects4J [9] offers reproducible Java bugs. However, these are often limited in scope, with fewer projects and less metadata, such as duplicate links or resolution details.
Our proposed dataset, GitBugs, addresses these limitations by aggregating bug reports from diverse platforms (GitHub, Jira, and Bugzilla) across multiple projects. By offering datasets of varying sizes per repository, it supports flexible training setups—ranging from smaller datasets suitable for current cost-constrained large language model applications to larger datasets ideal for training traditional machine learning models at scale.
II-B Applications in Research
Bug report datasets have supported a variety of research efforts, including duplicate detection [10, 11], automated triaging [12], and severity prediction [13]. The growing application of natural language processing (NLP) and deep learning in software engineering [14] has amplified the need for large and small, high-quality datasets.
GitBugs enables such applications by offering realistic, multi-domain bug data with detailed annotations. It supports reproducible experimentation across traditional machine learning and large language model (LLM)-based techniques.
III Dataset Description
III-A Data Sources
The GitBugs dataset aggregates bug reports from nine well-established open-source software projects: Cassandra, Firefox, Hadoop, HBase, Mozilla Core, VS Code, Seamonkey, Spark, and Thunderbird. These projects span multiple domains including distributed systems, browsers, IDEs, and cloud infrastructure. We sourced the reports from these bug trackers using project-specific APIs and scraping utilities.
III-B Data Collection Methodology
We gathered bug reports using a combination of RESTful API calls (for Jira-based systems like Apache projects) and HTML scraping or CSV archive parsing (for Bugzilla-based systems like Firefox and Thunderbird). We filtered reports to exclude non-bug entries such as feature requests or tasks using available metadata fields such as issue type.
III-C Data Volume and Statistics
Table I summarizes the dataset at the project level, reporting the total number of bug submissions and the percentage of duplicates. Mozilla Core contains the most significant number of reports, exceeding 85,000, followed by VS Code and Firefox. In contrast, projects like SeaMonkey and Cassandra have comparatively smaller datasets. Duplicate rates show considerable variation: VS Code and Thunderbird both exceed 25%, suggesting room for improvement in triage and reporting workflows.
Additional insights not shown in figures are outlined below:
  • Resolution time exhibits a heavy-tailed distribution: While some bugs are resolved within days, others remain open for months or years.
  • Bug priority correlates with outcomes: High-priority issues are more frequently resolved as Fixed or Won’t Fix, while low-priority ones are often closed as Duplicate or left unresolved.
  • Report lengths are highly variable: The median bug description is under 600 characters, though some exceed several thousand characters due to verbose logs or stack traces.
TABLE I: Summary statistics across projects in GitBugs.
Project Total Reports Duplicates Duplicate Rate (%)
Cassandra 4,612 300 6.5
Firefox 28,824 6,255 21.7
Hadoop 2,503 128 5.1
HBase 5,403 108 2.0
Mozilla Core 85,673 17,899 20.9
VS Code 32,829 9,272 28.2
Seamonkey 1,076 120 11.2
Spark 20,275 497 2.5
Thunderbird 15,192 4,200 27.6
Figure 1 illustrates monthly bug reporting trends. Mozilla Core consistently leads in report volume. Firefox shows prominent spikes in mid-2021 and 2024. VS Code and Thunderbird report at a steady, moderate pace, while Cassandra and SeaMonkey remain relatively low.
Figure 1: Monthly bug report trends from 2020 to 2024 across multiple projects. Mozilla Core consistently reports the highest volume, while other projects show lower and more variable trends over time.
Figures 2 and 3 compare bug resolution efficiency across projects. Spark shows the fastest resolution times, peaking at under 10 days. In contrast, SeaMonkey and Mozilla Core exhibit longer tails, reflecting extended issue lifetimes. Box plots reveal wide variance in resolution times for projects like Thunderbird and Mozilla Core, whereas VS Code and Spark demonstrate more consistent and efficient resolution behavior.
Figure 2: Kernel density estimates of bug resolution times across multiple projects. Spark shows the fastest resolution, while SeaMonkey has the longest tail, indicating slower bug fixes.
Figure 3: Distribution of bug resolution times across projects using box plots. Most projects exhibit a right-skewed distribution with many outliers; Mozilla Core, SeaMonkey, and Thunderbird show notably longer resolution times.
III-D Comparison table with existing datasets:
TABLE II: Comparison of Software Bug Report Datasets
Dataset Name Year Source Projects / Domain Size / Scope Key Features
PROMISE Repository[8] 2007 Various software projects Empirical software engineering data Benchmark datasets for software engineering research
Eclipse & Mozilla Defect Tracking Dataset[1] 2013 Eclipse (JDT, CDT, Platform, GEF), Mozilla (Firefox, Thunderbird, Bugzilla, Core) $\sim$215,000 bug reports Full bug lifecycle with update history; XML format
Defects4J[9] 2014 Java projects Real faults in Java programs Enables controlled testing studies
GitHub Issues Dataset (8M)[15] 2017 GitHub 8 million issues Large-scale dataset for issue tracking and analysis
Bughub[2] 2018 GitHub Free-text bug reports Designed for duplicate issue identification
BugRepo[16] 2018 Various software projects Dataset for duplicate bug report detection Supports research on duplicate bug report detection
Apache Bug-Fix Commits Dataset[6] 2019 55 Apache projects 10 years of bug-fix commits Links bug reports to commits; useful for bug-fix analysis
BugHunter[4] 2020 Various open-source Java projects Code elements (files, classes, methods) with bug information Automatically constructed bug dataset with code metrics; supports bug prediction research
Bugl[5] 2020 Cross-language (Java, Python, etc.) Cross-language bug reports Facilitates bug localization across languages
AndroR2[3] 2021 Android apps Manually reproduced bug reports Supports reproducibility studies
Eclipse Bugzilla Dataset[17] 2022 Eclipse Bug reports from Eclipse’s Bugzilla Structured dataset for bug analysis
Public Jira Dataset[18] 2022 Various projects using Jira Publicly available Jira issues Facilitates research on issue tracking systems
RegMiner[7] 2022 Various code repositories Replicable regression bug dataset Focuses on regression bugs
GitBugs (This Paper) 2025 Bugzilla, JIRA, Github 150K+ Bug Reports with mapped duplicates Duplicate Issue Identification, Reterival Augmented Generation, Prediction etc.
IV Potential Applications
The GitBugs dataset enables various research and industrial applications in software engineering, particularly bug report analysis, natural language processing, and machine learning.
IV-A Research Opportunities
GitBugs supports several high-impact research directions, including:
  • Duplicate bug detection: The availability of labeled duplicates enables training and evaluation of IR-based, graph-based, and neural models for identifying semantically similar bug reports.
  • Bug triaging: Supervised models can be developed to predict suitable developers, components, or priorities for new bug reports based on historical assignments and metadata.
  • Resolution and severity prediction: The structured fields Resolution, Priority, and Status provide valuable ground truth for classification tasks involving severity estimation and outcome forecasting.
  • Time-to-fix modeling: With timestamp data for creation and resolution dates, regression models can be trained to predict the expected resolution time of a bug.
  • Temporal and linguistic evolution: Longitudinal analysis of reporting patterns, language use, and resolution behaviors across projects and time can yield insights into software project dynamics and process maturity.
IV-B Industry Use Cases
Beyond academic settings, the dataset can be leveraged by software organizations for:
  • Tool benchmarking: Internal bug analysis tools (e.g., triage recommenders or deduplication engines) can be validated using standardized, labeled data from GitBugs.
  • LLM fine-tuning: The bug summaries and descriptions offer a rich corpus for pretraining or instruction tuning large language models for software maintenance tasks.
  • Custom QA pipelines: Real-world bug metadata and distributions help design and evaluate risk scoring, prioritization, and anomaly detection models tailored to in-house development workflows.
  • Training and education: GitBugs can be a hands-on resource in software engineering courses, enabling students to explore realistic bug data and apply analysis techniques.
IV-C Case Study: Trend Analysis and Prediction with GitBugs
We conducted a case study on Apache Cassandra bug reports to demonstrate the utility of the GitBugs dataset, reserving the final six months of data for testing. The study evaluated various machine learning (ML) models across key software engineering tasks: bug volume forecasting, severity classification, time-to-fix prediction, and trend detection.
Prediction Tasks.
Bug Volume Forecasting. We forecasted future bug report volumes using ARIMA and Prophet models trained on monthly historical data, evaluating the last six months reports. As shown in Figure 4, ARIMA produced a stable forecast, while Prophet reflected a sharper recent increase. Based on Mean Absolute Error (MAE), ARIMA outperformed Prophet with scores of 10.92 and 20.10, respectively, underscoring a trade-off between stability and sensitivity to recent trends.
Figure 4: Monthly bug report forecasts using ARIMA and Prophet models. Actual data (blue) is shown alongside ARIMA (orange) and Prophet (green) forecasts for 2024–2025.
Bug Priority Classification. We trained a multi—class classifier to predict High, Low, Normal, and Urgent bug priorities. As shown in Figure 5, it achieved strong performance on the dominant Normal class (F1-score: 0.90), while performance on minority classes was poor. Despite an overall accuracy of 82%, the macro-averaged F1-score was only 0.35, indicating significant class imbalance and the need for mitigation strategies such as re-sampling or cost-sensitive learning.
Figure 5: Confusion matrix for bug severity classification on the Cassandra dataset. Most samples are correctly classified as class 2, with moderate confusion between adjacent classes.
Time-to-Fix Prediction. We trained a regression model to predict the time-to-fix for bug reports. As illustrated in Figure 6, the predicted values often diverge from the actual values, particularly for reports with longer resolution times. The model yielded a Mean Absolute Error (MAE) of 86.06, a Root Mean Squared Error (RMSE) of 158.44, and an R2 score of –0.09, reflecting poor predictive performance. The negative R2 score indicates that the model underperforms even a naive predictor that always returns the mean. These results highlight the need for more informative features or alternative modeling approaches. The use of MAE as a performance metric is consistent with prior work [19].
Figure 6: Scatter plot of actual vs. predicted bug fix times. The red dashed line represents the ideal prediction (y = x). Most predictions underestimate the actual time to fix, indicating model bias toward shorter durations.
Trend Analysis.
Keyword Trends. We tracked domain-specific terms such as repair, auth, and timeout. Spikes in terms like repair and read correlated with maintenance-focused development phases, revealing shifts in team priorities.
Topic Modeling. Applying Latent Dirichlet Allocation (LDA) uncovered five dominant topics across the dataset: indexing, test automation, configuration cleanup, CVE/security, and flaky test failures. We illustrated this temporal evolution in Figure 7.
Figure 7: Monthly distribution of bug reports across LDA-derived topics from 2020 to 2024. Each topic shows fluctuating trends, with occasional spikes indicating bursts of activity in specific thematic areas.
Seasonal Decomposition. We applied STL decomposition to isolate monthly bug volume’s seasonal and trend components. The seasonal pattern aligned with quarterly release cycles, while the trend revealed a long-term decline in bug reports—suggesting increased system maturity, as shown in Figure 8.
Figure 8: STL decomposition of monthly bug reports: trend, seasonality, and residuals.
Duplicate Detection. To evaluate the dataset’s support for information retrieval tasks, we conducted a duplicate bug detection experiment using cosine similarity over bug summaries encoded with Sentence-BERT embeddings. For each of the 300 randomly selected query bugs, We retrieve the top 10 most similar candidates based on cosine similarity scores.
As illustrated in Figure 9, most top-10 similarity scores fell below 0.5, indicating that many duplicates remain linguistically subtle. Despite this challenge, the approach achieved a Recall@10 of 0.61, suggesting that semantic embedding techniques can retrieve a majority of known duplicates with high precision in the candidate set.
Figure 9: Histogram of cosine similarity scores between bug report embeddings. Most similarities fall below 0.5, indicating substantial semantic variation across reports, with a small cluster of highly similar cases near 1.0.
Retrieval-Augmented Generation.
Bug Explanation via Retrieval. To demonstrate GitBugs’ support for natural language workflows, we implemented a Retrieval-Augmented Generation (RAG) pipeline. Given a new bug report, the system retrieves semantically similar historical reports and conditions a language model to produce enriched explanations.
Input Bug Report:
When I press login on the settings page, nothing happens in Firefox Nightly.
Retrieved Examples:
  • Login button in about:preferences does not trigger any action when clicked.
  • Firefox login screen fails silently due to a missing DOM event listener on macOS.
  • Click events not captured on some buttons after upgrading to Nightly 120.0.
Generated Explanation:
This issue likely stems from a missing or broken event listener tied to the login button. In similar cases, the root cause was a regression in the event handling system introduced in a recent update. Suggested next step: check the associated click handler in the latest Nightly build and verify DOM readiness before event binding.
This example highlights how GitBugs enables language models to reason with contextual evidence, improving interpretability and alignment with prior developer behavior. Unlike zero-shot generation, RAG leverages historical grounding to support triage, summarization, and root cause hypothesis in a transparent manner.
Insights. This case study highlights GitBugs’ versatility in supporting various analytical tasks. It enables both predictive modeling and exploratory analysis of real-world development patterns. The findings also underscore several challenges, including class imbalance, high variance in fix durations, and limitations of textual signals—providing directions for future research on robust, domain-aware ML methods in software engineering.
V Conclusion and Future Work
V-A Summary
GitBugs represents a scalable, diverse, and richly annotated corpus of real-world software bug reports. By aggregating data from nine widely used open-source projects and unifying their structure, the dataset provides a valuable benchmark for empirical studies and machine learning research in software engineering.
With comprehensive fields such as Summary, Description, Status, Priority, and Resolution, the dataset supports a variety of research tasks including duplicate detection, triaging, severity prediction, and resolution time estimation. We believe GitBugs can facilitate reproducible experimentation and foster software analytics and innovation in tool development.

### v3

@@
-GitBugs: Bug Reports for Duplicate Detection, Retrieval Augmented Generation, and More
+GitBugs: Bug Reports for Duplicate Detection, Retrieval Augmented Generation, Triage, and More
 Abstract
@@
 IV Potential Applications
-The GitBugs dataset enables various research and industrial applications in software engineering, particularly bug report analysis, natural language processing, and machine learning.
+The GitBugs dataset supports a wide range of research and industrial applications in software engineering, particularly in bug report analysis, natural language processing, and machine learning.
+Table III presents a brief survey of existing studies on bug report analysis.
+TABLE III: Survey of traditional bug management tasks and emerging LLM-enabled expansions.
+Applications Representative Studies
+Traditional Tasks
+Duplicate Detection [62, 23, 38, 19, 56, 31, 35, 24, 20, 12, 18, 64]
+Severity Prediction [5, 65, 2, 8, 54, 57, 36, 34, 44, 1, 15]
+Component Assignment [42, 33, 11, 52, 58]
+Triage Automation [4, 63, 59, 60, 16, 28]
+LLM-Enabled Expansions
+Report Rewriting & Clarification [3, 9]
+Missing Information Inference [13, 41]
+Automated Reproduction Step Extraction [26, 27, 14, 51, 53, 61]
+Conversational Debugging [13, 47, 7]
 IV-A Research Opportunities